OpenStudy (anonymous):

A blueprint for a house states that 2.5 inches equals 10 feet. If the length of a wall is 12 feet, how long is the wall in the blueprint? a. 1 in. c. 3 in. b. 2 in. d. 4 in.

OpenStudy (anonymous):

another proportion....only this time, you have to set it up.. 2.5/10 = x/12 (2.5 in to 10 ft = x in to 12 ft.) now cross multiply

OpenStudy (anonymous):

ohkay. 2.5x=120

OpenStudy (anonymous):

120/2.5=48?